OpenStudy (anonymous):

the price of gasoline at a particular station is 1.5 euros per liter. An american student can use 33 euros to buy gasoline. knowing that 4 quarts make a gallon ant that 1 liter is close to 1 quart, she quickly reasons that she can buy how many gallons ofof gasoline?

OpenStudy (anonymous):

I have a feeling I've helped with this one before. But let's do it. If you can buy 1 litre with 1.5 euros how many litres can you buy with 33 euros is the first part. How do you get on with that part?

OpenStudy (anonymous):

4 qts makes a gallon and 1 liter is close to a quart, therefore, 4 liters make a gallon. Therefore, a gallon = (4 x 1.5) = 6 euros. And 33/6 = 5.5 gallons. or... 1.5/1 = 33/x (cross multiply) 1.5x = 33 x = 22 liters. (33 euro's buys 22 liters) And since 4 quarts make a gallon, and 1 liter is close to a quart, then 4 liters makes a gallon. 4 liters/1 gallon = 22 liters/x gallons 4x = 22 x = 5.5 gallons